# Slow-phonon control of spin Edelstein effect in Rashba $d$-wave altermagnets

### v1

@@ Abstract
Altermagnets have zero net magnetization yet feature spin-split bands that host spin-polarized states. Here, we investigate how slow lattice vibrations (phonons) influence both the intrinsic and externally induced spin polarizations in two-dimensional $d$-wave altermagnets. For the induced spin polarizations, we employ a Rashba continuum model with electron-phonon coupling (EPC) treated at the static-Holstein level and analyze the spin Edelstein effect using the Kubo linear-response formalism. We find that moderate-to-strong EPC progressively suppresses the induced polarization via both intraband and interband channels, with a critical coupling marking the onset of complete spin Edelstein depolarization. The depolarization transition arises from a phonon-induced energy renormalization that pushes the spin-split bands anisotropically above the chemical potential, leading to a complete collapse of the Fermi surface. While (de)polarization can occur even in the Rashba non-altermagnetic phase, it remains isotropic. The presence of altermagnetism makes it anisotropic and breaks the conventional antisymmetry between spin susceptibilities that occurs with pure spin-orbit coupling, rendering the effect highly relevant for spintronic applications. We further investigate how the phonon coupling to the altermagnetic order, Rashba spin-orbit strength, and carrier doping collectively tune the depolarization transition. Our findings demonstrate that phonon scattering (e.g., through various substrates) offers a powerful means for on-demand control of spin polarization, enabling reversible switching between spin-polarized and depolarized states—a key functionality for advancing spin logic architectures and optimizing next-generation spintronic devices.
@@ I Introduction
Although altermagnets offer exciting opportunities for spintronic applications, the impact of lattice dynamics (phonons) on their intrinsic and extrinsic spin polarizations is still largely unexplored. Electron-phonon coupling (EPC) can induce quasiparticle renormalization and enable phonon-assisted transport processes in both electronic and magnetic systems [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Very recently, Irosh [58] demonstrated that electron pairing mediated by dispersive phonons in altermagnets can give rise to reentrant superconductivity as a function of temperature. Leraand et al. [59] further showed that, within the weak EPC regime, the dominant superconducting instability in altermagnets is odd in momentum and even in spin, hosting spin-polarized Cooper pairs. In addition, Steward et al. [60] reported that coupling the altermagnetic order to lattice vibrations leads to the emergence of a hybridized paramagnon-polaron mode. In another study, Hodt et al. [61] found that electron-phonon scattering processes can strongly enhance the spin conductivity of altermagnets at finite frequencies. He et al. [62] have also emphasized the pivotal role of electron-phonon scattering in shaping both the efficiency and the directionality of charge-to-spin conversion in altermagnets. Nonetheless, despite phonon’s ubiquity in electronic systems, the role of slow phonons in shaping the low-energy spin excitations and spin Edelstein effect in Rashba altermagnets has so far received little attention and remains entirely uncharted.
@@
In typical two-dimensional (2D) Rashba systems, the electronic bandwidth far exceeds the characteristic phonon energy scale [63, 64], implying that phonons evolve on much slower timescales than electronic degrees of freedom. This separation of timescales forms the basis of many theoretical approaches to spin-related phenomena in altermagnets, which remain largely unexplored. Building on the above recent developments in understanding EPC effects in altermagnets, this work addresses the open question of how slow phonons, in conjunction with intrinsic spin polarization, affect the induced spin Edelstein polarizations in $d$-wave altermagnets. To capture this interplay, we construct a minimal continuum model at low temperatures that includes both RSOC and linear EPC (between lattice displacements and both spin-degenerate and spin-split states) at the static-Holstein level. Using the Kubo formalism, we compute induced spin Edelstein polarization over an extensive range of parameters, revealing distinctive fingerprints of the combined effects of lattice dynamics and the underlying crystalline symmetry in altermagnets.
@@
Our analysis uncovers several key insights: (i) EPC induces a net magnetization, or equivalently an out-of-plane spin polarization, and renormalizes the intrinsic spin splitting in $d$-wave altermagnets; (ii) intraband and interband scattering processes strongly suppress the Edelstein response with increasing EPC, eventually leading to a critical depolarization point where the induced spin polarization vanishes permanently; (iii) the presence of altermagnetism enhances anisotropy and modifies the usual antisymmetry between spin susceptibilities, producing directionally dependent (de)polarization; and (iv) Rashba strength and doping provide tunable knobs that shift the onset of depolarization and govern whether spin polarization is stabilized.
@@
Our findings establish phonon engineering as a promising route to optimize next-generation spintronic devices. Enhancing phonon scattering (e.g., using various substrates which produce different EPCs) can suppress nonequilibrium spin accumulation, reducing spin-related dissipation in spin-orbit torque devices. Moreover, on-demand switching between spin-polarized and depolarized states—crucial for spin logic architectures where spin rather than charge encodes information [65]—relies on controlled depolarization to reset logic elements, erase stored information, and prevent spin leakage.
@@
Figure 1: (a) Real-space representation of the staggered spin configuration on a square lattice of a pristine $d_{x^{2}-y^{2}}$-wave altermagnet with phonons included and without gating, where the spin-dependent orbital texture alternates between sublattices. Sublattice states are linearly coupled to slow Holstein phonons, depicted as shaded (vibrating) sites. (b) Corresponding band structure along high-symmetry paths near the $\Gamma$-point, displaying spin-split dispersions for spin-up and spin-down channels.
@@ II.1 Pristine $d_{x^{2}-y^{2}}$-wave altermagnet
The low-energy quasiparticle excitations of a 2D Rashba-coupled $d_{x^{2}-y^{2}}$-wave altermagnet, Fig. 1(a), can be described by an effective two-band Hamiltonian acting in spin space. It consists of three main contributions: (i) a kinetic energy term, proportional to the identity matrix $\sigma_{0}$, which is spin-independent; (ii) an anisotropic spin-splitting term, proportional to $\sigma_{z}$, that originates from the $d$-wave altermagnetic order; and (iii) a RSOC term, proportional to $\sigma_{x}$ and $\sigma_{y}$, which breaks inversion symmetry, stemming from the gate electrodes [41, 42]. Explicitly, the pristine Rashba Hamiltonian reads
@@
$\alpha_{\vec{k}}=\frac{\hbar^{2}}{2m}\,(k_{x}^{2}+k_{y}^{2})\quad{\rm and}\quad\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{2m}\,(k_{x}^{2}-k_{y}^{2})\,$, (2)
with $m$ the effective electron mass and $0<\beta<1$ parametrizing the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting. In this model, $\beta_{\vec{k}}$ even in $\vec{k}$ ($\pi/a<k_{x}<\pi/a$ and $\pi/a<k_{y}<\pi/a$ with lattice constant $a$) preserves $C_{4}$ rotational symmetry, while RSOC term breaks inversion symmetry, thus, spin locks to momentum. Accordingly, RSOC wants to twist spin around the Fermi surface, while altermagnetism distorts it anisotropically. We briefly note that although our present study focuses on the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting, other symmetry patterns can be readily incorporated by rotating the plane of the altermagnet. For instance, a $45^{\circ}$ rotation of the plane yields a $d_{xy}$-wave altermagnet characterized by $\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{m}\,k_{x}k_{y}$.
Diagonalization of Eq. (1) yields the band dispersions
@@ II.2 Phonon-dressed $d_{x^{2}-y^{2}}$-wave altermagnet
where $Q_{i}$ represents the displacement of the ion at site $i$, $P_{i}$ is the conjugate momentum, $M$ is the ion mass, and $\omega_{\rm p}$ is the characteristic phonon frequency.
In altermagnets at half filling, the spin-sublattice symmetry ensures that the electronic density is spatially uniform, thus, any static lattice distortion would break this symmetry without lowering the energy. To maintain this symmetry, we impose that the net force on the lattice vanishes at half filling. Concretely, this means that the phonon displacement $Q_{i}$ couples to density fluctuations rather than the absolute electron density, so that $\langle n_{i,\uparrow}+n_{i,\downarrow}\rangle=1$. This centers the lattice equilibrium configuration at half filling and ensures that the electron-phonon interaction vanishes in the absence of charge or spin-density fluctuations. This is a standard assumption in electron-phonon models [68, 69, 70, 71, 72, 73, 74] and corresponds to the standard local linear coupling. However, the intrinsic altermagnetic order of the sublattices allows us to in general incorporate spin-dependent coupling constants, so that the electron-phonon interaction is described by
$\mathcal{H}_{\text{e-p}}=\sum_{i}\left[g_{\uparrow}\left(n_{i,\uparrow}-\tfrac{1}{2}\right)+g_{\downarrow}\left(n_{i,\downarrow}-\tfrac{1}{2}\right)\right]Q_{i}\,.$ (6)
We stress that in our model, we subtract the half filling density $1/2$ rather than the instantaneous average density. In standard electron-phonon models with a quadratic phonon potential, this choice would only shift the equilibrium lattice configuration and could be absorbed into the chemical potential without changing the physics. However, in our case, the fixed chemical potential renders the choice physically meaningful: it guarantees that the coupling term vanishes at half filling where the lattice symmetry remains unbroken.
Crucially, since lattice displacements $Q_{i}$ are even under time reversal while $n_{i,\uparrow}-n_{i,\downarrow}$ is odd, the bilinear $(n_{i,\uparrow}-n_{i,\downarrow})Q_{i}$ is time-reversal-odd and therefore can appear only in a system that already breaks time-reversal symmetry. Although the displacement vectors vary in direction across different sites and spin states, their overall pattern follows a staggered arrangement: spin-up along $x$ experiences $Q_{i}$ while spin-down experiences $Q_{i}$, with the roles reversed along $y$. Accordingly, a nonzero $g_{\uparrow/\downarrow}$ is symmetry-allowed in altermagnets, but it is forbidden in time-reversal-symmetric crystals unless an external time-reversal-odd field is present. In fact, lattice distortions alter bond lengths and angles, and in altermagnets with intrinsically anisotropic spin textures, such geometric modulations generically induce spin-asymmetric corrections to the electronic structure, which at low energies project onto a local spin-dependent potential.
We therefore introduce symmetric and antisymmetric combinations of the coupling constants as
$g=\frac{g_{\uparrow}+g_{\downarrow}}{2},\quad\tilde{g}=\frac{g_{\uparrow}-g_{\downarrow}}{2}\,.$ (7)
We define the spin-dependent couplings as $g_{\uparrow}=g+\tilde{g}$ and $g_{\downarrow}=g-\tilde{g}$. Substituting these into the general EPC term in Eq. (6), we obtain
@@
This expression naturally separates into two contributions: a symmetric part coupling to the total local electron density $n_{i}=n_{i,\uparrow}+n_{i,\downarrow}$ and an antisymmetric part coupling to the local spin imbalance $m_{i}=n_{i,\uparrow}-n_{i,\downarrow}$. As previously noted, lattice displacements $Q_{i}$ modulate crystal fields and exchange/hopping in the presence of RSOC, producing a spin-asymmetric on-site potential. In a two-sublattice altermagnet the net magnetization induced by $\tilde{g}$ becomes finite for staggered (or otherwise symmetry-breaking) displacement patterns, yielding a macroscopic piezomagnetic response. Thus, microscopically, $\tilde{g}$ directly maps onto the piezomagnetic coefficient measurable in displaced altermagnets [75, 76, 77, 78, 60]. We emphasize that in the absence of altermagnetism, the effective coupling $\tilde{g}$ is forbidden, implying that the two parameters $\beta$ and $\tilde{g}$ are interdependent.
@@
The phonon-induced terms $gQ_{0}$ and $\tilde{g}Q_{0}$ play distinct roles in the low-energy electronic structure. The shift $gQ_{0}$ renormalizes the chemical potential and, away from half filling, provides a route for phonon-assisted tuning of the carrier density. In contrast, the spin-dependent term $\tilde{g}Q_{0}$ acts as a dynamical Zeeman-like field, lifting the spin degeneracy and enabling transient spin polarization or spin-current control. When $\tilde{g}=0$, only the $gQ_{0}$ term is present, producing an overall band shift without modifying the spin texture arising from $\beta_{\vec{k}}$ or RSOC; the spectrum therefore remains isotropic along the $\Gamma\!\to\!X$ and $\Gamma\!\to\!Y$ directions. By contrast, a finite $\tilde{g}Q_{0}$ introduces opposite shifts for spin-up and spin-down bands, leading to spin-dependent anisotropies even though the coupling itself is momentum independent.
@@
$M\omega_{\rm p}^{2}Q_{0}={}-g\,[n_{\rm e}-1]-\tilde{g}\,m\,$, (10)
where $n_{\rm e}=n_{\uparrow}+n_{\downarrow}$ is the total density and $m=n_{\uparrow}-n_{\downarrow}$ is the imbalance density, with
@@
$m={}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k},s}f_{\vec{k},s}\langle\tilde{\psi}_{\vec{k},s}|\sigma_{z}|\tilde{\psi}_{\vec{k},s}\rangle\,$, (11b)

### v2

@@ Abstract
Altermagnets have zero net magnetization yet feature spin-split bands. Here, we investigate how slow lattice vibrations (phonons) influence both the intrinsic and externally induced spin polarizations in two-dimensional $d$-wave altermagnets. For the induced spin polarization, we employ a Rashba continuum model with electron-phonon coupling (EPC) treated at the static-Holstein level and analyze the spin Edelstein effect using the Kubo linear-response formalism to probe EPC-induced contributions. We find that, under a specific symmetry-lowering pattern like a piezomagnetically active strain that explicitly breaks the inherent $C_{4}\mathcal{T}$ symmetry, moderate-to-strong EPC progressively suppresses the induced polarization via both intraband and interband channels, with a threshold coupling marking the onset of complete spin Edelstein depolarization. The depolarization arises from a phonon-induced energy renormalization that leads to a complete collapse of the Fermi surface. While (de)polarization can occur even in the Rashba non-altermagnetic phase, it remains isotropic. The presence of altermagnetism makes it anisotropic and breaks the conventional antisymmetry between spin susceptibilities that occurs with pure spin-orbit coupling, rendering the effect highly relevant for spintronic applications. We further investigate how the phonon coupling to the altermagnetic order, Rashba spin-orbit strength, and carrier doping collectively tune the depolarization. Our findings demonstrate that static phononic effects offer a powerful means for on-demand control of spin polarization, enabling reversible switching between spin-polarized and depolarized states—a key functionality for advancing spin logic architectures and optimizing next-generation spintronic devices.
@@ I Introduction
Although altermagnets offer exciting opportunities for spintronic applications, the impact of lattice vibrations (phonons) on their intrinsic and extrinsic spin polarizations has not yet been fully understood. Electron-phonon coupling (EPC) can renormalize quasiparticles and enable phonon-assisted transport in electronic and magnetic systems [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Very recently, Iorsh [58] showed that dispersive phonons can mediate reentrant superconductivity in altermagnets. Leraand et al. [59] demonstrated that, in the weak EPC regime, the leading superconducting instability is odd in momentum and even in spin, yielding spin-polarized Cooper pairs. Steward et al. [60] reported hybrid paramagnon-polaron modes arising from coupling between altermagnetic order and phonons. Moreover, Hodt et al. [61] found a strong phonon-induced enhancement of finite-frequency spin conductivity, while He et al. [62] highlighted the role of electron-phonon scattering in controlling charge-to-spin conversion.
@@
In typical two-dimensional (2D) Rashba systems, the electronic bandwidth far exceeds the characteristic phonon energy scale [63, 64], implying that phonons evolve much more slowly than electronic degrees of freedom. This separation of timescales underlies many theoretical treatments of spin phenomena in altermagnets, yet remains largely unexplored. Despite the ubiquity of phonons, the influence of slow phonons (low frequency/long period)—common in real materials—on low-energy spin excitations and the spin Edelstein effect in Rashba altermagnets remains largely unexplored. By taking a step back, this approach simplifies the theory and clarifies phenomena that are difficult to extract from more complex models. Motivated by recent advances on EPC effects in altermagnets, we address how slow phonons, together with intrinsic spin polarization, influence the spin Edelstein response in $d$-wave altermagnets. To this end, we develop a minimal low-temperature continuum model incorporating RSOC and linear EPC (between lattice displacements and both spin-degenerate and spin-split states) at the static-Holstein level. Using the Kubo formalism, we compute the induced spin Edelstein polarization across a wide parameter range, uncovering clear signatures of the interplay between slow phonons and crystalline symmetry.
@@
Our analysis uncovers several key insights: (i) EPC induces a net magnetization, or equivalently an out-of-plane spin polarization, and renormalizes the intrinsic spin splitting in $d$-wave altermagnets; (ii) intraband and interband transitions strongly suppress the Edelstein response with increasing EPC, eventually leading to a depolarization, where the induced spin polarization vanishes permanently; (iii) the presence of altermagnetism enhances anisotropy and modifies the usual antisymmetry between spin susceptibilities, producing directionally dependent (de)polarization; and (iv) Rashba strength and doping provide tunable knobs that shift the onset of depolarization and govern whether spin polarization is stabilized. We emphasize that phonon-induced effects are equilibrium band-structure phenomena, independent of the Edelstein effect, which is a nonequilibrium transport response. Our goal is not to invoke an additional microscopic depolarization mechanism beyond this band-structure renormalization, but to highlight that the Edelstein effect serves as a sensitive probe of the phonon-renormalized Fermi surface.
We note that a static lattice distortion, stemming from a piezomagnetically-active strain pattern, induces a uniform spin polarization (magnetization) by breaking pre-existing $C_{4}\mathcal{T}$ symmetry (a combined action of time-reversal symmetry and the lattice rotation or mirror symmetry) [65, 66, 67, 68, 69, 70, 71, 66, 72]), allowing lattice distortions to convert staggered spin textures into a uniform spin imbalance.
@@
Within a fixed-chemical-potential ensemble, the Edelstein susceptibility vanishes when distortion-induced band renormalization removes all electronic states from the Fermi level. The disappearance of the response therefore reflects the collapse of the Fermi surface rather than a breakdown of spin-momentum locking at fixed carrier density. From this perspective, the Edelstein effect should be viewed as a probe of the underlying Fermi-surface topology. Our findings thus establish phonon engineering as a promising route to optimize next-generation spintronic devices. Moreover, on-demand switching between spin-polarized and depolarized states—crucial for spin logic architectures where spin rather than charge encodes information [73]—relies on controlled depolarization to reset logic elements, erase stored information, and prevent spin leakage.
@@
Figure 1: (a) Real-space representation of the staggered spin configuration on a square lattice of a pristine $d_{x^{2}-y^{2}}$-wave altermagnet with phonons included and without gating, where the spin-dependent orbital texture alternates between sublattices. The out-of-plane arrows denote two symmetry-related sublattices, $A$ and $B$, which host electrons with opposite spin polarization. Sublattice states are linearly coupled to slow Holstein phonons, depicted as shaded (vibrating) sites. Two types of EPC are shown: the conventional symmetric coupling $g$ for each sublattice/spin and the staggered antisymmetric coupling $\tilde{g}$ arising from $C_{4}\mathcal{T}$ symmetry breaking. (b) Corresponding band structure along high-symmetry paths near the $\Gamma$-point, displaying spin-split dispersions for spin-up and spin-down channels.
@@ II.1 Pristine $d_{x^{2}-y^{2}}$-wave altermagnet
The low-energy quasiparticle excitations of a 2D Rashba-coupled $d_{x^{2}-y^{2}}$-wave altermagnet can be described by an effective two-band Hamiltonian acting in spin space. It consists of three main contributions: (i) a kinetic energy term, proportional to the identity matrix $\sigma_{0}$, which is spin-independent; (ii) an anisotropic spin-splitting term, proportional to $\sigma_{z}$, that originates from the $d$-wave altermagnetic order; and (iii) a RSOC term, proportional to $\sigma_{x}$ and $\sigma_{y}$, which breaks inversion symmetry, stemming from the gate electrodes [41, 42, 74]. The arrows in Fig. 1(a) denote two symmetry-related sublattices, labeled $A$ and $B$, which host electrons with opposite out-of-plane spin polarizations. Explicitly, the pristine Rashba Hamiltonian reads
@@
$\alpha_{\vec{k}}=\frac{\hbar^{2}}{2m_{\rm e}}\,(k_{x}^{2}+k_{y}^{2})\quad{\rm and}\quad\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{2m_{\rm e}}\,(k_{x}^{2}-k_{y}^{2})\,$, (2)
with $m_{\rm e}$ the effective electron mass and $0<\beta<1$ parametrizing the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting. In this model, $\beta_{\vec{k}}$ even in $\vec{k}$ ($\pi/a<k_{x}<\pi/a$ and $\pi/a<k_{y}<\pi/a$ with lattice constant $a$) preserves $C_{4}$ rotational symmetry, while RSOC term breaks inversion symmetry, thus, spin locks to momentum. Accordingly, RSOC wants to twist spin around the Fermi surface, while altermagnetism distorts it anisotropically. We briefly note that although our present study focuses on the $d_{x^{2}-y^{2}}$-wave anisotropy of the spin splitting, other symmetry patterns can be readily incorporated by rotating the plane of the altermagnet. For instance, a $45^{\circ}$ rotation of the plane yields a $d_{xy}$-wave altermagnet characterized by $\beta_{\vec{k}}=\frac{\hbar^{2}\beta}{m_{\rm e}}\,k_{x}k_{y}$.
Diagonalization of Eq. (1) yields the bare band dispersions
@@
The electronic Hamiltonian of an altermagnet is invariant under the combined symmetry $C_{4}\mathcal{T}$. Consequently, the band energies must satisfy $E_{\vec{k}}(\sigma)=E_{(C_{4}\mathcal{T})\,\vec{k}}\!\left((C_{4}\mathcal{T})\,\sigma\right)$. At the $\Gamma$ point, $\vec{k}=0$ is invariant under $C_{4}$ rotations, and time reversal $\mathcal{T}$ also leaves $\vec{k}=0$ unchanged. However, $\mathcal{T}$ flips the spin, $\sigma\rightarrow-\sigma$. Therefore, applying $C_{4}\mathcal{T}$ at $\Gamma$ yields $E_{0}(\sigma)=E_{0}(-\sigma)$, which guarantees spin degeneracy at the $\Gamma$ point. Thus, as long as $C_{4}\mathcal{T}$ symmetry is preserved, no spin splitting can occur at $\vec{k}=0$.
In a 2D altermagnet, spin polarization is locked to the sublattice degree of freedom rather than producing a net magnetization. The unit cell consists of two symmetry-related sublattices, $A$ and $B$, which host electrons with opposite spin polarization. Specifically, sublattice $A$ predominantly supports spin-up electrons, while sublattice $B$ supports spin-down electrons, such that the local spin expectation values satisfy $\langle\vec{S}_{A}\rangle=-\langle\vec{S}_{B}\rangle$ and the total magnetization vanishes. Despite the absence of macroscopic magnetization, the lack of combined $\mathcal{PT}$ symmetry allows for momentum-dependent spin splitting of electronic bands even in the absence of spin-orbit coupling. As a result, the Bloch states exhibit a sublattice-resolved spin texture, leading to characteristic altermagnetic spin-momentum locking protected by crystal symmetries rather than relativistic effects.
@@ II.2 Phonon-dressed $d_{x^{2}-y^{2}}$-wave altermagnet
where $Q_{i}$ represents the displacement of the ion within the unit cell $i$, $P_{i}$ is the conjugate momentum, $M$ is the ion mass, and $\omega_{\rm p}$ is the characteristic phonon frequency.
The general linear EPC at the lattice level reads
$H_{\mathrm{e\text{-}p}}=\sum_{i,\alpha,\mu}\psi_{i}^{\dagger}g_{\alpha}^{\mu}\psi_{i}\,d_{i\alpha}^{\mu},$ (6)
where $\alpha$ labels atoms within the unit cell, $d_{i\alpha}^{\mu}$ atomic displacements, and $g_{\alpha}^{\mu}$ is a single-particle operator acting in the orbital/sublattice space, describing how electrons couple to a displacement (or force) of atom $\alpha$ along direction $\mu$. Expanding in scalar phonon normal modes $Q^{\nu}$ with polarization vectors $e_{\alpha}^{(\nu)\mu}$ gives $d_{i\alpha}^{\mu}=\sum_{\nu}e_{\alpha}^{(\nu)\mu}Q^{\nu}_{i}$ and $H_{\mathrm{e\text{-}p}}=\sum_{i,\nu}Q^{\nu}_{i}\psi_{i}^{\dagger}\Big[\sum_{\alpha,\mu}g_{\alpha}^{\mu}e_{\alpha}^{(\nu)\mu}\Big]\psi_{i}.$ Projecting onto a $s$-band subspace, $\psi_{i}=\sum_{s}c_{i}^{s}|u_{s}\rangle$, where $c^{s}_{i}$ is the fermionic annihilation operator for an electron in band $s$ localized at unit cell $i$, encoding the amplitude/operator degree of freedom associated with all atomic orbitals within unit cell $i$ that contribute to band $s$, we define
$(\hat{O}_{\nu})_{s^{\prime}s}\equiv\sum_{\alpha,\mu}\langle u_{s^{\prime}}|g_{\alpha}^{\mu}|u_{s}\rangle\,e_{\alpha}^{(\nu)\mu},$ (7)
so that the Hamiltonian becomes $H_{\mathrm{e\text{-}p}}=\sum_{i,\nu}Q^{\nu}_{i}\,c^{\dagger}_{i}\hat{O}_{\nu}c_{i}$. The matrix element $\langle u_{s^{\prime}}|g_{\alpha}^{\mu}|u_{s}\rangle$ is the band-resolved coupling vertex associated with moving atom $\alpha$ along direction $\mu$, giving rise to the corresponding intra- and interband coupling amplitudes. All directional and atomic information enters only parametrically via $\hat{O}_{\nu}$. Different lattice displacements therefore correspond simply to different values of these coefficients.
With adiabatic phonons ($\nu=0$) and following the spinors in Eq. (4) for our low-energy model in the presence of static lattice distortion, as well as Eq. (1) with $\lambda_{\rm R}=0$, $\hat{O}_{0}$ decomposes as (see Appendix A for a detailed derivation)
$\hat{O}_{0}=\frac{g_{A}^{x}+g_{B}^{y}}{2}\sigma_{0}+\frac{g_{A}^{x}-g_{B}^{y}}{2}\sigma_{z}\,.$ (8)
This is a minimal model as a proof-of-principle, demonstrating how the pristine Rashba altermagnet can be modified by a static lattice distortion. The $\sigma_{0}$ term corresponds to a sublattice-symmetric deformation potential, while the $\sigma_{z}$ term represents a staggered EPC that breaks sublattice symmetry and can generate a Zeeman-like contribution in the altermagnetic phase. In the pristine altermagnet, the degeneracy at the $\Gamma$ point is protected by the combined action of time-reversal symmetry and the lattice rotation (or mirror) symmetry that exchanges the two sublattices, i.e., by the $C_{4}\mathcal{T}$ symmetry. To have the $\sigma_{z}$ term, we propose a static lattice distortion to break $C_{4}\mathcal{T}$ symmetry. This distortion drives atomic displacements within the unit cell, thereby lowering the symmetry. This can occur in three distinct ways: by breaking the $C_{4}$ rotational symmetry, by breaking time-reversal ($\mathcal{T}$) symmetry, or by breaking both $C_{4}$ and $\mathcal{T}$ such that their combined operation is no longer a symmetry. In all cases, the protecting $C_{4}\mathcal{T}$ symmetry is no longer present and an an out-of-plane Zeeman-like term of the form $\propto\sigma_{z}$ can be generated. To only break $C_{4}$, one can think of structural anisotropy that lowers rotation symmetry and this is physically possible through uniaxial strain, distortion, or substrate-induced anisotropy [65, 66, 67, 68, 69, 70, 71, 66, 72]. To only break $\mathcal{T}$, one can add a Zeeman magnetic field to destroy the altermagnetic protection [77]. To break both $C_{4}$ and $\mathcal{T}$, one can apply a circularly polarized light (Floquet), which breaks $\mathcal{T}$ and mirror symmetries [78, 79, 80].
In practice, a physically realistic mechanism must involve a specific symmetry-lowering distortion pattern that explicitly breaks the relevant $C_{4}\mathcal{T}$ symmetry. Since we involve a deformation process by considering phonons, breaking $C_{4}$ is the best solution. We identify a piezomagnetically active strain [65, 66, 67, 68, 69, 70, 71, 66, 72] as a plausible microscopic origin of the static lattice distortion considered here. This is both theoretically well-justified and experimentally relevant. However, a detailed analysis of the underlying symmetry-breaking mechanism—such as that discussed in Ref. [72]—goes beyond the scope of the present minimal model and will be addressed in future work.
By setting $Q_{i}^{0}=Q_{i}$, the distortion-induced electron–phonon Hamiltonian can be written as
$\mathcal{H}_{\text{e-p}}=\sum_{i}\left[g^{x}_{A}\left(n_{i,A}-\tfrac{1}{2}\right)+g^{y}_{B}\left(n_{i,B}-\tfrac{1}{2}\right)\right]Q_{i}\,.$ (9)
The shaded sites in Fig. 1(a) schematically represent a staggered phonon mode that couples to the electronic density. The subtraction of $1/2$ from the electron number ensures that the phonon coordinate $Q_{i}$ couples only to deviations from half-filling, so that $Q_{i}=0$ corresponds to the undistorted equilibrium lattice [81, 82, 83, 84, 85]. Thus, the EPC is nonzero only when the electron density deviates from $1/2$, leading to a finite phonon displacement. Since $g_{A}^{x}$ acts only on sublattice $A$ with spin-up and $g_{B}^{y}$ only on sublattice $B$ with spin-down, we choose $g_{A}^{x}=g_{\uparrow}$ and $g_{B}^{y}=g_{\downarrow}$ across the system. On the other hand, since we do not perform ab initio calculations to determine direction-dependent EPC constants for different displacement patterns, we introduce the symmetric and antisymmetric combinations (see Fig. 1(a))
$g=\frac{g_{\uparrow}+g_{\downarrow}}{2},\quad\tilde{g}=\frac{g_{\uparrow}-g_{\downarrow}}{2}\,$, (10)
in which we define the spin-dependent couplings as $g_{\uparrow}=g+\tilde{g}$ and $g_{\downarrow}=g-\tilde{g}$. In the absence of precise values for $g_{\uparrow}^{x}$ and $g_{\downarrow}^{y}$, sweeping a broad parameter range for $g$ and $\tilde{g}$ effectively spans the relevant lattice-displacement directions.
Substituting these into the general EPC term in Eq. (9), we obtain
@@
This expression naturally separates into two contributions: a symmetric part coupling to the total local electron density $n_{i}=n_{i,\uparrow}+n_{i,\downarrow}$ and an antisymmetric part coupling to the $m_{{\rm s},i}=n_{i,\uparrow}-n_{i,\downarrow}$, the difference between electronic densities on the two altermagnetic sublattices. In our model, the bilinear term $Q\,(n_{\uparrow}-n_{\downarrow})$ does not represent a local spin on a single site, but the staggered, sublattice-resolved electron density in the altermagnet. Since the symmetry-lowering piezomagnetically-active strain patterns break $C_{4}\mathcal{T}$ symmetry [65, 66, 67, 68, 69, 70, 71, 66, 72], this coupling is allowed. Therefore, no compensating frequency is required, and the coupling is fully consistent with the symmetry of the altermagnetic Hamiltonian.
@@
The phonon-induced terms $gQ_{0}$ and $\tilde{g}Q_{0}$ play distinct roles in the low-energy electronic structure. The shift $gQ_{0}$ renormalizes the chemical potential and provides a route for phonon-assisted tuning of the carrier density. In contrast, the spin-dependent term $\tilde{g}Q_{0}$ acts as a Zeeman-like field, lifting the spin degeneracy and enabling transient spin polarization or spin-current control. When $\tilde{g}=0$, the spectrum therefore remains isotropic along the $\Gamma\!\to\!X$ and $\Gamma\!\to\!Y$ directions. By contrast, a finite $\tilde{g}$ introduces opposite shifts for spin-up and spin-down bands, leading to spin-dependent anisotropies even though the coupling itself is momentum independent.
While the present model is formulated at a minimal, proof-of-principle level, the breaking of $C_{4}\mathcal{T}$ symmetry induced by a static lattice distortion remains operative in realistic multi-sublattice systems [4, 86]. In particular, if a staggered field emerges from the interplay between magnetic and nonmagnetic sublattices mediated by lattice distortions in realistic materials, this still validates the underlying symmetry-based mechanism discussed here. Although material-specific details may renormalize the effective electronic-structure parameters and the strength of the EPCs, the fundamental distortion-induced phenomena described here remain symmetry-allowed and robust.
@@
$M\omega_{\rm p}^{2}Q_{0}={}-g\,[n_{\rm e}-1]-\tilde{g}\,m_{\rm s}\,$, (13)
where $n_{\rm e}=n_{\uparrow}+n_{\downarrow}$ is the total density and $m_{\rm s}=n_{\uparrow}-n_{\downarrow}$ is the imbalance density, with
@@
$m_{\rm s}={}$ $\frac{1}{N_{\vec{k}}}\sum_{\vec{k},s}f_{\vec{k},s}\langle\tilde{\psi}_{\vec{k},s}|\sigma_{z}|\tilde{\psi}_{\vec{k},s}\rangle\,$, (14b)